OpenStudy (joehoffy):

can someone teach me how to find X in here: x^3 = 125

OpenStudy (triciaal):

Do you know how to find the prime factors of a number?

OpenStudy (joehoffy):

no

OpenStudy (triciaal):

Do you know what a prime number is?

OpenStudy (joehoffy):

like sqrt you mean? i know that, and i know cube root but i just want to know how to do it on a calculator

OpenStudy (triciaal):

A prime number is a number that can only be divided by 1 and itself

OpenStudy (triciaal):

2 is the only even prime number

OpenStudy (triciaal):

Here we want the factors of 125

OpenStudy (triciaal):

Prime numbers 1,3,5,7,11,13,17,19,23,29 etc What is the lowest that can divide into 125 ?

OpenStudy (triciaal):

Do you see that 125 = 5*5*5?

OpenStudy (triciaal):

X^3=5^3

OpenStudy (joehoffy):

yes

OpenStudy (joehoffy):

so its 5^3

OpenStudy (triciaal):

This is simple enough to see x = 5. But to continue the process on the left you have x^3 so to get x you would take the cube root For an equation whatever you do to one side also do to the other so take the cube root of 125 which is 5.

OpenStudy (triciaal):

I showed the prime factorization so you can do without a calculator. You would group by 2's for square root, 3's for cube root etc

OpenStudy (triciaal):

Any questions?

OpenStudy (joehoffy):

all i ever wanted to know is how to find cube root in my calc

OpenStudy (triciaal):

Cube root is to the power 1/3

OpenStudy (joehoffy):

so it would have been 1/3 of 125?

OpenStudy (triciaal):

3*(1/3) = 1

OpenStudy (joehoffy):

calculated industries scientific calculator

OpenStudy (triciaal):

No not 1/3 of 125 It would be 125 raised to the power 1/3

OpenStudy (joehoffy):

no

OpenStudy (triciaal):

put 125^(1/3)=

OpenStudy (joehoffy):

41.6

OpenStudy (triciaal):

The ^ is most of the time a 2nd function key

OpenStudy (joehoffy):

there is one. that is like X^?

OpenStudy (joehoffy):

it gives me 41.666666666666666666666666666667

OpenStudy (joehoffy):

I DID IT GAVE ME 41.666666666666666666666666666667

OpenStudy (triciaal):

Use the parenthesis or put 0.33. Order of operations is still important

OpenStudy (joehoffy):

OOH IT WORKED NOW i got 5

OpenStudy (triciaal):

You are welcome

OpenStudy (joehoffy):

thank you